Wyslij-po Manual

version 3.3, 8 February 2021

Sergey Poznyakoff.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007-2021 Sergey Poznyakoff Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “Wyslij-po Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this Manual, like GNU software. Help the software be free.”

Table of Contents

1 Introduction to Wyslij-po . . . . . . . . . . . . . . . . . . . . . . 1 2 Submitting Translations. . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Verification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Sending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 Debugging, Verbose Output and Dry-run

Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4 Configuration Files.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Language Tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 How to Report a Bug . . . . . . . . . . . . . . . . . . . . . . . . . 13 Appendix A Downloading wyslij-po . . . . . . . . . . . 15 Appendix B How to Build the Package . . . . . . . 17 Appendix C Invocation Summary.. . . . . . . . . . . . . 19 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

1 Introduction to Wyslij-po

It is a common practice for a program to be documented and to produce its diagnostics in English. However, the waste majority of us would prefer their computer to speak our native language instead. This goal is already achieved for many free software programs, and is going to be achieved for still more of them using the methods called internationalization and localization. These terms are often written abbreviated as i18n and l10n, correspondingly. In short, i18n consists in modifying a program in such a way that, while still producing its output in English, it would allow this output to be translated to any given language on the fly. Similarly, l10n consists in supplying an already internationalized program enough information for it to be able to translate its output to a particular language. Both terms are discussed in further detail in Section “I18n” in GNU gettext manual.

While internationalization is usually taken care of by programmers, localization is the responsibility of translators. Usually this process consists of creating a special file, called PO file, which is basically a table of all the messages the program ever issues, which, for each such message, lists its translation to a given language (see Section “PO Files” in GNU gettext manual).

When translator finishes writing a PO file, he submits it via electronic mail to the free Translation Project (http://translationproject.org), or TP for short, which acts as a repository for the existing translations. It is from the Translation Project that maintainers download PO files for inclusion into their projects, when preparing them for the release.

There are several ways to submit a translation. First, it can, of course, be submitted manually. Secondly, you can use the M (po-send-mail) command in GNU Emacs po-mode (see Section “PO Files” in GNU gettext manual). Finally, you can use some specialized program to do that.

If you choose the last possibility, wyslij-po is for you. It is a specialized program which verifies PO files and submits them to the TP.

The program name^{1} means ‘send-po’ in Polish and is pronounced roughly like

‘vishleey-po’.

The program is built using GNU Mailutils (http://www.gnu.org/software/ mailutils) — a powerful collection of libraries and utilities for handling electronic mail.

1 It should more properly have been written ‘wyślij-po’.

2 Submitting Translations.

To submit one or several PO files, simply give them as arguments to wyslij-po in the command line, like this:

 $ wyslij-po pl.po uk.po bg.po

For each file, specified in the command line, the utility will do the following steps:

1. Retrieve PO headers PO headers contain important information about the translation. These headers are stored in the first entry of each PO file.

2. Restore canonical PO name.

A canonical PO name consists of the project name and version, separated by dash, which are followed by a dot, the language code as per ISO 639, and the suffix ‘.po’. To determine the canonical name, wyslij-po uses the contents of Project-Id-Version and Language PO headers. The value of Project-Id-Version supplies the name and version of the project this file is for, and the value of Language gives the ISO-639 abbreviation of the language name. For example, if the two headers in question were:

     Project-Id-Version: cflow 1.2
     Language: pl

then the canonical name for that file is cflow-1.2.pl.po. If the PO file lacks the Language header, wyslij-po analyzes the Language-Team header. The full language name extracted from this header is looked up in a language table file to find the corresponding ISO-639 code (see Chapter 5 [Langtab], page 11).

3. Verify the consistency of PO file. This step ensures that the file corresponds to the latest version of the package in question and that it does not contain obvious formatting errors. For the detailed description of the checks performed, see Section 2.1 [Verification], page 3, below.

4. Form the message to be submitted. The message is a MIME 1.0 message, whose body contains the compressed PO file, and the headers contain the necessary registration information (see Section 2.2 [Message], page 4).

5. Send the message.

See Section 2.3 [Sending], page 6, below.

The following sections describe each step in detail.

2.1 Verification

Before constructing a submission message, wyslij-po first verifies if the PO file in question matches several important conditions. It is done in order to reduce the unwanted traffic and the impact on the TP mail server and robot.

As of version 3.3, the following tests are implemented:

1. PO revision date.

The values of the PO headers PO-Revision-Date and POT-Creation-Date are com-

pared. If POT-Creation-Date is newer than PO-Revision-Date, the PO file is obviously outdated and therefore it is not submitted.

This test is controlled by the following command line options:

--verify-po-time

Verify PO modification time (default).

--no-verify-po-time

Disable PO modification time verification.

In configuration file, this test is controlled by the verify-po-time boolean statement (see Chapter 4 [Config], page 9).

2. Package version This test verifies if this PO file is created for the latest version of the package available. To verify this the following algorithm is used. First, the textual domain url is built using the following schema:

     http://translationproject.org/domain/${domain}.html

where ${domain} is replaced with the project name obtained from the Project-Id-Version PO header. Then, the contents of this page is retrieved, and the following extended regular expression is applied to it:

     The current template for this domain is <a href=".[^"]*/(.[^"]*)">

If the expression matches, the value of the first parenthesized subexpression is retrieved, the optional ‘.pot’ suffix is removed from it and the result is compared with the value of Project-Id-Version PO header, converted as described in [canonical-name], page 3. The two strings must be equal in order for the test to succeed. This test is controlled by the following command line options:

--verify-version

Enable package version verification (default).

--no-verify-version

Disable package version verification.

In configuration file, this test is controlled by the verify-version boolean statement (see Chapter 4 [Config], page 9). Two configuration file statements are provided for tuning this test. The tp-url statement changes the textual domain template (see Chapter 4 [Config], page 9) and the pot-regex statement changes the regular expression (see Chapter 4 [Config], page 9).

2.2 Message

The minimal message headers are:

 From: sender
 To: robot-address
 Subject: TP-Robot canonical-name
 X-Mailer: wyslij-po 3.3

Where:

sender The email of the sender. Unless otherwise specified, it is constructed by concatenating the actual user name, a ‘@’ sign, and the host name of the local host.

This value can be set explicitly using --from (-F) command line option. When used without additional arguments, as in:

     $ wyslij-po -F pl.po

this option instructs wyslij-po to extract sender name from the value of Last-Translator PO header. This is equivalent to setting from-po true in the configuration file (see Chapter 4 [Config], page 9). This option can also be used with an explicit argument, which specifies the email to be used as a sender address. It is important to remember that when the short option form is used, this argument must follow the option letter immediately, without any intervening whitespace. If the long option form is used, this argument must be delimited from the option name by an equals sign with no extra whitespace on either side of it. The following invocations illustrate this:

     wyslij-po -Ftrans@gnu.org.ua
     wyslij-po --from=trans@gnu.org.ua

This can be requested in configuration file by using from statement (see Chapter 4 [Config], page 9).

robot-address

The email of the TP robot. By default it is:

     robot@translationproject.org

It can be changed using --to (-t) command line option, or to configuration statement (see Chapter 4 [Config], page 9).

canonical-name The canonical name of the PO file. It is determined as described in [canonicalname], page 3.

Additionally, you can require the message to be copied to any number of emails by using --cc command line option or cc configuration statement. In both cases, the argument is a string containing a comma-separated list of valid email addresses, as in:

$ wyslij-po --cc "tp-keeper@dom.ain,root@remote.org"

If you wish to make those additional addresses invisible, use the --bcc command line option (or bcc configuration statement) instead.

The --fcc option specifies the folder to store the copy of the message into. Its argument is any mailbox URL supported by Mailutils (see Section “URL” in GNU Mailutils Manual). For example, the following option will place copies of the PO files in the MH folder

Mail/sent:
$ wyslij-po --fcc mh://Mail/sent

In configuration file, this is done via the fcc statement (see Chapter 4 [Config], page 9). The PO file contents is by default sent in compressed form. To produce it, gzip program

is used (see Section “Overview” in Gzip Manual). Notice, that the gzip binary must be in your PATH, otherwise wyslij-po won’t be able to find it. You can specify the full path to the binary and/or supply any additional options by using the --gzip (--compress) command line option. For example, to use /usr/bin/gzip with the highest possible compression level, use:

     $ wyslij-po --gzip "/usr/bin/gzip -c --best"

Use the compress statement to set the compression command from the configuration file (see Chapter 4 [Config], page 9).

Notice, that when setting compression command (either from command line or from configuration file), you must always include -c option in the gzip command line.

2.3 Sending

As any mailutils-based program, wyslij-po uses a special logical entity called mailer in order to actually submit messages. By default it uses the MTA facilities of the machine it runs on, so you usually don’t have to explicitly specify which mailer to use. There may, however, be some cases when this becomes necessary. This section describes how to do so.

A mailer is specified by its URL. An URL begins with a protocol specification. Two protocol specifications are currently supported: ‘sendmail’ and ‘smtp’. The former means to use a sendmail-compatible program to send mails. Such a program must be able to read mail from its standard input and must support the following options:

-oi Do not treat ’.’ as message terminator.

-f addr Use addr as the address of the sender.

-t Get recipient addresses from the message.

These conditions are met by most existing MTA programs, such as exim or postfix (to say nothing of sendmail itself).

Following the protocol specification is the mailer location, which is separated from it with a colon. For the ‘sendmail’ protocol, the mailer location sets the full file name of the sendmail-compatible MTA binary, for example:

     sendmail:/usr/sbin/sendmail

A special form of sendmail URL, consisting of protocol specification only (‘sendmail:’) is also allowed. It means “use the sendmail binary from the _PATH_SENDMAIL macro in your /usr/include/paths.h file”. This is the default mailer.

The ‘smtp’ protocol means to use an SMTP server directly. In this case the mailer location consists of two slashes, followed by the IP address or host name of the SMTP server, and, optionally, the port number. If the port number is present, it is separated from the rest of URL by a colon. For example:

     smtp://remote.server.net
     smtp://remote.server.net:24

The URL of the mailer to use is set using --mailer (-M) command line option. For example:

     $ wyslij-po --mailer=smtp://127.0.0.1 pl.po

An alternative way is to set it in the configuration file, using mailer statement (see Section “Mailer Statement” in GNU Mailutils Manual).

As an example of a rather extreme approach, here is how to deliver mails directly to the TP mail server:

     $ wyslij-po --mailer=smtp://mail.vrijschrift.org ...

3 Debugging, Verbose Output and Dry-run Mode.

Usually wyslij-po operates silently. It issues a diagnostics only if something serious has happened (an error, for example). You may wish to change this behavior in order to trace the progress of the submission.

The --verbose command line option instructs the program to increase its verbosity level. Each occurrence of this option in the command line raises the level by one, so to get the verbosity level 3 you would type:

$ wyslij-po -vvv ...

The bigger the level, the more information will be produced on the output.

Here is an example of the output produced in level 1:

$ wyslij-po -v pl.po
pl.po: cflow-1.2.pl.po
Message headers:
Content-transfer-encoding: base64
Content-Type: application/x-gzip; name="cflow-1.2.pl.po.gz"
MIME-Version: 1.0
From: trans@gnu.org.ua
To: robot@translationproject.org
Subject: TP-Robot cflow-1.2.pl.po
X-Mailer: wyslij-po 3.3

Here, the first line shows the input file name and the corresponding canonical PO name for it. The remaining lines show RFC 2822 message headers.

The verbosity level 2 shows additional information about the progress:

$ wyslij-po -vv pl.po
Sourcing language table `./wyslij-po.lc'
Restoring canonical file name for `pl.po'
Verifying POT file version for `mailutils'
Verifying creation vs. revision date
pl.po: cflow-1.2.pl.po
executing gzip -c pl.po...
Message headers:
Content-transfer-encoding: base64
Content-Type: application/x-gzip; name="cflow-1.2.pl.po.gz"
MIME-Version: 1.0
From: trans@gnu.org.ua
To: robot@translationproject.org
Subject: TP-Robot cflow-1.2.pl.po
X-Mailer: wyslij-po 3.3
Submitting pl.po

The levels 3 and 4 print lots of information concerning actual SMTP transaction. These are mainly useful for debugging.

It is often desirable to know what wyslij-po would do, without actually sending translations. The special command line option --dry-run (-n) is designed for this purpose. When this option is given, wyslij-po will do all its usual processing, except for actually sending the result. This option automatically sets verbose level to 1 (of course you can push it higher by supplying additional -v options along with --dry-run):

$ wyslij-po --dry-run pl.po
pl.po: cflow-1.2.pl.po
Message headers:
Content-transfer-encoding: base64
Content-Type: application/x-gzip; name="cflow-1.2.pl.po.gz"
MIME-Version: 1.0
From: trans@gnu.org.ua
To: robot@translationproject.org
Subject: TP-Robot cflow-1.2.pl.po
X-Mailer: wyslij-po 3.3
NOT submitting

4 Configuration Files.

If you have a set of commonly used wyslij-po options, it is convenient to keep them in a configuration file, instead of typing them again each time you invoke the program. Wyslijpo uses the Mailutils concept of configuration files, which is described in detail in Section “configuration” in GNU Mailutils Manual.

There are two possible locations for the wyslij-po configuration. System-wide configuration is kept in the main configuration file (see Section “configuration” in GNU Mailutils Manual), in section program wyslij-po, or in the file named wyslij-po in the configuration include directory, if the include statement is used (see Section “Include” in GNU Mailutils Manual). Per-user configuration is kept in file ~/.wyslij-po.

To set default mailer (similar to --mailer option, see Section 2.3 [Sending], page 6), use the mailer statement, See Section “Mailer Statement” in GNU Mailutils Manual.

[Configuration] compress command

Set command line of the utility used for file compression. For example:

     compress "/usr/bin/bzip2 -c";

See [gzip], page 5, for a detailed discussion of file compression. See also Appendix C [Invoke], page 19, for another way to set this value.

[Configuration] langtab file

Set the name of the language table file. See Chapter 5 [Langtab], page 11.

[Configuration] to email

Set TP robot (recipient) email address. See Section 2.3 [Sending], page 6.

[Configuration] cc email

Set emails of carbon copy recipients. Email may contain multiple email addresses, separated by commas:

     cc "<tp@localhost>,<bitbucket@remote.tld>";

Notice the use of quote characters. See also Appendix C [Invoke], page 19, for another way to set this value.

[Configuration] bcc email

Set emails of carbon copy recipients. See also Appendix C [Invoke], page 19, for another way to set this value.

[Configuration] fcc folder

Store a copy of the sent message in the specified mail folder. Folder can be any valid mailbox URL (see Section “URL” in GNU Mailutils Manual). The following example instructs wyslij-po to keep copies of the messages it sends in a MH folder ‘+po-out’:

     fcc "mh:~/Mail/po-out";

[Configuration] from email

Set sender e-mail address. See Section 2.2 [Message], page 4, for a detailed description. See also Appendix C [Invoke], page 19, for another way to set this value.

[Configuration] from-po bool

Read sender e-mail address from the value of Last-Translator PO header. It is equivalent to specifying the --from (-F) command line option without argument.

[Configuration] verify-po-time bool

Controls whether to verify PO modification time before sending the message. See Section 2.1 [Verification], page 3, for a description of this test.

[Configuration] verify-version string

Controls whether to verify package version before sending the message. See Section 2.1 [Verification], page 3, for a description of this test.

[Configuration] tp-url template

Sets template for the textual domain URL. See Section 2.1 [Verification], page 3, for a description of its purpose. Template must be a valid HTTP URL. It must contain at least one occurrence of ${domain}, which will be replaced by the actual name of the textual domain.

[Configuration] pot-regex regex

Set the regular expression matching the name of the last POT file for the textual domain. See Section 2.1 [Verification], page 3, for a detailed description of its purposes. Regex must be an extended regular expression (see Section “Extended regular expressions” in GNU sed) with exactly one parenthesized group.

5 Language Tables.

To recreate the canonical PO name (see [canonical-name], page 3) for a given translation, wyslij-po needs to know the language code associated with this language. Normally this information is kept in the Language header of the PO file^{1}. If this header is not present, wyslij-po will first obtain the full language name from the Language-Team PO header, and then use this value as a look-up key in a set of language tables. If the key is found, its value gives the language code.

This language code lookup is retained as a fallback mechanism for handling old PO files and will be removed in future releases. This chapter describes the format of language table files used by wyslij-po.

The language tables are stored in plain text files. Their format is as follows:

• Empty lines are ignored.

• Any lines whose first non-whitespace character is ‘#’ are comments and are ignored as well.

• Unless ignored by the above reasons, any line is split at the whitespace. The first word in line gives the language code (usually corresponding to ISO 639). The rest of words is taken as a literal string that gives the full language name (the look-up key) for that code.

For example, the following entry:

     pt_BR Brazilian Portuguese

means that when the string ‘Brazilian Portuguese’ is encountered at the beginning of the Language-Team PO header, the language code ‘pt_BR’ will be used.

There are two language table files: a system-wide and a per-user one.

The system-wide language table is stored in file wyslij-po.lc in your system configuration directory (by default it is /usr/local/etc, but can be changed at compile time).

The per-user language table is read after the system-wide one. It is located in the user’s home directory and is named .wyslij-po.lc. You can also supply another file name, using

the --langtab command line option, e.g. --langtab=mytab.lc.

The contents of both files are concatenated and stored in a single table.

Notice that it is not an error if a language table file does not exist. Wyslij-po will silently ignore non-existing files. Therefore, if you get the similar error:

     wyslij-po: Unknown language: Bulgarian

you should make sure the right tables are read, by running wyslij-po with the -vv (or, better yet, --dry-run --verbose) command line option. In normal case, it will at least read the site-wide language table, which will be reflected in its output:

     $ wyslij-po --dry-run -v bg.po
     Sourcing language table `/usr/local/etc/wyslij-po.lc'
     Restoring canonical file name for `bg.po'

If it does not, make sure you install the table or supply its alternate name using -- langtab. Otherwise, inspect the table and add an entry for the language in question.

1 The Language header is present in PO files created by gettext version 0.18 or later.

The package is shipped with a system wide configuration file built using the data from http://translationproject.org/team, so most of the time you won’t have to edit it (unless, of course, new language teams are added to the TP)

6 How to Report a Bug

Email bug reports to bug-wyslij-po@gnu.org.ua.

As the purpose of bug reporting is to improve software, please be sure to include maximum information when reporting a bug. The information needed is:

• Version of the package you are using.

• Compilation options used when configuring the package.

• Conditions under which the bug appears.

Appendix A Downloading wyslij-po

Stable releases of wyslij-po can be downloaded from http://download.gnu.org.ua/
pub/release/wyslij-po^{1}.

Each tarball in this directory is GPG signed and the corresponding signature is available in file with the .sig suffix. To verify the tarball, be sure to download both the tarball and its signature (e.g., for the latest version: wyslij-po-3.3.tar.gz and wyslij-po-3.3.tar.gz.sig ). Then run a command like this:

     gpg --verify wyslij-po-3.3.tar.gz.sig

If that command fails because you don’t the author’s public key, then run this command to import it:

     gpg --keyserver keys.gnupg.net --recv-keys 3602B07F55D0C732

A symbolic link to the latest release (and its signature) is provided: http://download.

gnu.org.ua/pub/release/wyslij-po/wyslij-po-latest.tar.gz.

Pre-release and alpha versions of the package can be downloaded from http://

download.gnu.org.ua/pub/alpha/wyslij-po.

Finally, development of the wyslij-po can be tracked from https://puszcza.gnu.

org.ua/projects/wyslij-po.

1 Traditionally, the FTP protocol is supported as well:

ftp://download.gnu.org.ua/pub/release/wyslij-po

Appendix B How to Build the Package

Building wyslij-po is quite straightforward: 1.

Make sure you have installed the GNU Mailutils, version 3.4 or later. 2.

   Run ./configure

3.

   Run make

4.

   Run make install

Appendix C Invocation Summary.

This appendix lists all wyslij-po command line options. For each option are given a short description and a reference to the part of this manual that fully describes it.

--bcc=email-list

Set blind carbon-copy recipients (see Section 2.2 [Message], page 4).

--cc=email-list

Set carbon-copy recipients (see Section 2.2 [Message], page 4).

--compress=command
--gzip=command

Use this command for compression (see Section 2.2 [Message], page 4).

--dry-run

-n Do nothing, print what would have been done (see Chapter 3 [Dry-run], page 7).

--fcc=folder

Store copy of each message sent in the given folder.

--from[=email]

-Femail Set sender address (see Section 2.2 [Message], page 4).

--help

-? Give a short help summary

--langtab=file

Set the name of the per-user language table file (see Chapter 5 [Langtab], page 11).

--mailer=url

-M url Use the specified url as the default mailer. See Section 2.3 [Sending], page 6.

--no-verify-po-time

Disable PO modification time verification. See Section 2.1 [Verification], page 3.

--no-verify-version

Disable package version verification. See Section 2.1 [Verification], page 3.

--to=email

-t email Set TP robot (recipient) email address. See Section 2.3 [Sending], page 6.

--usage Produce a listing of all command line options.

--verbose

-v Produce verbose output. See Chapter 3 [Dry-run], page 7.

--verify-po-time

Verify PO modification time. See Section 2.1 [Verification], page 3.

--verify-version

Make sure the submitted PO file matches the latest package version. See Section 2.1 [Verification], page 3.

--version

-V Print program version

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles.

You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 29 30 Wyslij-po Manual

Concept Index

This is a general index of all issues discussed in this manual

B

bcc, --bcc option, described . . . . . . . . . . . . . . . . . . . 5 building wyslij-po . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

C

cc, --cc option, described . . . . . . . . . . . . . . . . . . . . . . 5 Command line options . . . . . . . . . . . . . . . . . . . . . . . . . 19 compress, --compress option, described . . . . . . . 5 compress, --compress option, summary . . . . . . 19 configuration file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

D

dry-run, --dry-run option, described. . . . . . . . . . 8 dry-run, --dry-run option, summary. . . . . . . . . 19

F

fcc, --fcc option, described . . . . . . . . . . . . . . . . . . . 5 fcc, --fcc option, summary . . . . . . . . . . . . . . . . . . 19 from, --from option, described . . . . . . . . . . . . . . . . 4 from, --from option, summary . . . . . . . . . . . . . . . 19

G

GNU Mailutils. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gzip, --gzip option, described . . . . . . . . . . . . . . . . 5 gzip, --gzip option, summary . . . . . . . . . . . . . . . 19

H

help, --help option, summary . . . . . . . . . . . . . . . 19

I

i18n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

l18n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 langtab, --langtab option, described . . . . . . . . 11 langtab, --langtab option, summary. . . . . . . . . 19 language code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 language tables, files . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 language tables, format . . . . . . . . . . . . . . . . . . . . . . . . 11 Localization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

M

mailer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 mailer URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 mailer, --mailer option, described . . . . . . . . . . . . 6 mailer, --mailer option, summary . . . . . . . . . . . 19 Mailutils . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

N

                                                no-verify-po-time, --no-verify-po-time

option, summary . . . . . . . . . . . . . . . . . . . . . . . . . 19

                                                no-verify-version, --no-verify-version

option, summary . . . . . . . . . . . . . . . . . . . . . . . . . 19

O

Options, command line . . . . . . . . . . . . . . . . . . . . . . . . 19

P

PO file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 PO headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PO verification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

S

sending messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

T

to, --to option, described . . . . . . . . . . . . . . . . . . . . . . 5 to, --to option, summary . . . . . . . . . . . . . . . . . . . . 19

TP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Translation Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

U

URL, mailer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 usage, --usage option, summary . . . . . . . . . . . . . 19

V

verbose, --verbose option, described. . . . . . . . . . 7 verbose, --verbose option, summary. . . . . . . . . 19 Verification of PO files . . . . . . . . . . . . . . . . . . . . . . . . . . 3

                                                verify-po-time, --verify-po-time option,

summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

                                                verify-version, --verify-version option,

summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 version, --version option, summary. . . . . . . . . 19

W

wyslij-po . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 wyslij-po, building . . . . . . . . . . . . . . . . . . . . . . . . . . . 17